# Reference-based Metrics Disprove Themselves in Question Generation

### v1

@@ 4.3 Human Evaluation
We recruit volunteer annotators, all fluent English speakers, to evaluate both model-generated questions and human-written questions. We use the 48 test instances from HotpotQA. For each instance, annotators evaluate four questions: RefQ, GPT-3.5 (zero-shot), BART-base, and AnnoQ. To ensure impartiality, we anonymize and shuffle these candidate questions before presenting them to the annotators. They are tasked with rating each question based on naturalness, answerability, and complexity, using a 3-point scale for each criterion (with 2 as acceptance and 0 as rejection). Additionally, we sum the individual scores together to calculate a combined score that reflects the question’s overall quality. We obtain three annotations per question and use the average of these as the standard for human judgment. The Pearson correlation coefficient between ratings given by our annotators are 0.67. The rating rubric is available in Appendix A.4.
@@
We calculate the Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients to measure the agreement between all metrics, including NACo, and human judgment, as reported in Table 3. This comparison considers correlation with both individual criteria and the overall question quality.
Table 3 reveals that NACo demonstrates the highest correlation with human evaluation for individual criteria in 7 out of 9 scores, and it ranks second-highest in the remaining 2 scores. Notably, NACo exhibits the strongest agreement with human judgment concerning the overall quality of questions across all correlation metrics, outperforming the state-of-the-art RQGUE metric by $6\%$ in both $r$, $\rho$, and $\tau$ correlations.
@@ 5 Related Work
Evaluation Metrics for Question Generation: The evaluation of QG models relies heavily on automatic evaluation, where examples that score well and poorly on the automatic metrics serve as indicators of potential issues that new approaches should address. The most commonly used metrics are reference-based metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), BERTScore Zhang et al. (2019), and BLEURT Sellam et al. (2020). Based on correlation with human judgment, there have been studies attempting to challenge the effectiveness of these reference-based metrics and propose reference-free evaluation mechanism for QG Nema and Khapra (2018); Ji et al. (2022); Mohammadshahi et al. (2023). Our study, however, questions the competency of reference-based metrics by replicating the data collection process of benchmarks and introducing new references.
@@ 6 Conclusion
In this work, we questioned the competency of reference-based metrics in providing an accurate assessment for question generation. We replicated the data collection process used for benchmark datasets, which allowed us to gather candidate questions qualified as new references. Our analysis highlights the shortcomings of reference-based metrics in differentiating new references from less desirable candidates, assigning significantly lower scores to the former. Even the recently introduced reference-free metric, RQUGE, face difficulties in this regard. To address these challenges, we introduce NACo, a multi-dimensional, reference-free metric designed to bridge the gap between automated evaluation and human judgment in question generation. Our experimental results showcase that NACo, leveraging the Chain-of-Thought capabilities of Large Language Models for question answering, not only meets the expectations for quantitative QG metrics but also achieves state-of-the-art alignment with human evaluation.
@@
A limitation of our work speaks to the required access to a reasonable number of references to assess domain-specific or dataset-specific complexity. Future works can investigate how to account for expected complexity in scenarios where references are limited and difficult to collect. Moreover, NACo, like other reference-free metrics for QG, is subject to the performance of the underlying QA model. Specifically, the constraints of GPT-3.5 in answering complex, multi-hop questions might have limited NACo’s ability to evaluate valid references closer to the upperbound. Future directions should explore evaluation frameworks that are robust to variations in QA model performance.
@@ A.1 What makes a good question?
Figure 4 illustrates the gap between current automatic QG metrics and human evaluated metrics, where two questions using similar words can have opposite qualities. This gap can be explained by the fact that existing automatic metrics do not directly address any of the criteria that human annotation often looks for in a question. To address this challenge, our metric integrates the human perspective of a "good" question: naturalness, answerability, and complexity, into the evaluation pipeline.
@@ A.2 Prompt for CoT-QA
      Step 1…
@@
      Step 2…
@@
A.3 Additional results for reference-based metrics
We report the performance of QG competitors across reference-based metrics, using questions created by our annotators (AnnoQ) as references, in Table 4 and 5.
QG Competitor Ref-based metrics Ref-free metrics Our metric
B B-RT R-L BSc Q-B QA-S R-Q Na. Ans. Cmp. NACo
LM-generated
BART-base 17.92 -0.30 43.29 92.03 37.24 -0.37 4.62 99.60 66.75 69.15 73.30
GPT-3.5 (few-shot) 21.04 -0.19 44.61 92.44 38.49 -0.37 4.56 100.00 65.23 69.63 73.67
BART-clue-RefQ 13.88 -0.37 40.43 91.74 33.2 -0.38 4.56 99.73 63.94 69.24 69.97
Human-validated
RefQ 12.78 -0.31 37.83 91.52 31.32 -0.38 4.89 99.87 68.67 68.93 75.09
AnnoQ 100.00 100.00 100.00 100.00 100.00 -0.37 4.71 99.73 67.70 70.33 74.01
AnnoQ-clue-RefQ 42.42 0.06 62.54 94.62 55.73 -0.38 4.76 99.87 66.34 70.36 74.21
Table 4: SQuAD - Performance of different QG methods on NACo and other existing metrics. The evaluation uses our annotated questions (AnnoQ) as references, with GPT-3.5 as the underlying QA system. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
QG Competitor Ref-based metrics Ref-free metrics Our metric
B B-RT R-L BSc Q-B QA-S R-Q Na. Ans. Cmp. NACo
LM-generated
BART-base 12.26 -0.87 30.11 87.05 24.26 -0.28 2.78 64.58 31.00 68.56 41.25
GPT-3.5 (zero-shot) 13.36 -0.84 33.34 88.94 20.46 -0.28 4.25 100.00 71.55 70.72 73.60
GPT-3.5 (few-shot) 14.19 -0.69 34.54 89.45 23.94 -0.27 4.21 99.31 67.31 72.64 71.09
BART-clue-RefQ 12.78 -0.87 34.02 88.78 27.02 -0.28 3.31 90.28 46.84 79.41 63.17
Human-validated
RefQ 13.64 -0.69 30.94 89.22 26.94 -0.27 4.21 99.31 59.62 83.02 74.33
AnnoQ 100.00 100.00 100.00 100.00 100.00 -0.27 4.32 100.00 76.11 82.05 83.43
AnnoQ-clue-RefQ 33.66 -0.35 50.60 93.08 43.93 -0.27 4.44 100.00 71.46 81.46 81.60
Table 5: HotpotQA - Performance of different QG methods on NACo and other existing metrics. The evaluation uses our annotated questions (AnnoQ) as references, with GPT-3.5 as the underlying QA system. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
A.4 Instructions for Human Evaluation
@@
A.5 Experiment Details
@@
For GPT3.5 (gpt3.5-turbo) QG and CoT-QA, we employ the default hyperparameters with a temperature of 0.7 and top_p of 1. In our CoT-QA experiments on SQuAD, given the larger sample size of 750 and cost constraints, we conducted a single run. For our CoT-QA experiments on HotpotQA, we carried out 3 runs on the 50-examples sample and report the average scores from the three responses.

### v2

@@ 4.1 Experimental Setup
Human Evaluation: We recruit volunteer annotators, all fluent English speakers, to evaluate both model-generated questions and human-written questions, using 48 test examples from HotpotQA. For each example, annotators evaluate four questions: RefQ, GPT-3.5 (zero-shot), BART-base, and AnnoQ, displayed in randomized and anonymized order. Evaluators rate each question based on naturalness, answerability, and complexity, using a 3-point scale for each criterion. Additionally, we sum the individual scores together to calculate a combined score that reflects the question’s overall quality. We obtain three annotations per question and use the average of these as the standard for human judgment. The Pearson correlation coefficient between ratings given by our annotators are 0.67. The rating rubric is available in Appx. A.4.1.
4.2 Results
Failure of reference-based metrics: We report QG competitors’ performance on various metrics, including NACo, using RefQ as the reference in Tbl. 1 for SQuAD. Even though RefQ, AnnoQ, and AnnoQ-clue-RefQ are all qualified as valid questions, reference-based metrics rate them with significant differences. In the SQuAD dataset, BLEU scores for RefQ, AnnoQ, and AnnoQ-clue-RefQ are 100, 12.78, and 27.43, respectively (Tbl. 1). However, NACo score these three groups questions similarly, with RefQ, AnnoQ, and AnnoQ-clue-RefQ scoring 75.09, 74.01, and 74.21, respectively (Tbl. 1). Similar patterns are observed in the HotpotQA and TedEdQA datasets, as detailed in Tbl. 6 and Tbl. 7.
According to reference-based metrics, models that learn from training data either through finetuning (like BART-base) or demonstration (like GPT-3.5) are scored significantly higher than our annotators, who lack access to the training data. For instance, in the case of SQuAD, BART-base is scored higher than AnnoQ by almost $7\%$ according to BLEU-4, reported in Tbl. 1. As reference-based metrics measure syntactic and semantic similarity, the use of a single reference can disqualify our annotated questions from being considered reference materials, resulting in a misleading portrayal of a valid group of candidate questions.
Effectiveness of NACo: Referring to our analysis of four groups of candidate questions for HotpotQA in Fig. 1, NACo uniquely succeeds in separating all four groups by significant margins, unlike the seven existing metrics. The newly collected multi-hop questions in Group 1, which satisfy all criteria for HotpotQA questions, achieve the highest average NACo score of $0.83$. They are followed by the questions in Group 2, lacking in complexity, with a score of $0.75$; Group 3, lacking in naturalness, with a score of $0.19$; and Group 4, lacking in answerability, with a score of $0.03$.
@@ 4.2 Results
We calculate the Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients to measure the agreement between all metrics, including NACo, and human judgment, as reported in Tbl. 2. This comparison considers correlation with both individual criteria and the overall question quality. Tbl. 2 reveals that NACo demonstrates the highest correlation with human evaluation for individual criteria in 7 out of 9 scores, and it ranks second-highest in the remaining 2 scores. Notably, NACo exhibits the strongest agreement with human judgment concerning the overall quality of questions across all correlation metrics. This observation is consistent across different underlying LLMs.
QG Competitor B R-Q NACo Human
LM-generated
BART-base 12.97 2.78 41.25 2.35
GPT-3.5 (zero-shot) 7.94 4.25 73.60 4.20
Human-validated
RefQ 100.00 4.21 74.33 4.77
AnnoQ 13.64 4.32 83.43 5.02
Table 3: HotpotQA - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original HotpotQA questions (RefQ) as references, with GPT-3.5 as the underlying QA system for NACo.
4.3 Analysis
NACo vs. Reference-based Metrics: Tbl. 3 indicates that reference-based metrics rate BART-base questions slightly lower than AnnoQ (by $0.67\%$ according to BLEU), whereas NACo shows a much larger gap ($42.18\%$). Upon manually reviewing the questions generated by BART-base, we noticed a considerable number of them were not actual questions but rather statements using similar wording to the reference question RefQ. This observation is validated by our human evaluators, detailed in A.4.2. Fig. 3 provides a case study where BERTScore, the reference-based metric most aligned with human judgment Mohammadshahi et al. (2023), favored BART-base generation over the human annotated question, even though the former was not formatted as a question. This incompetency of reference-based metric can be explained by the fact that BART, when finetuned on the HotpotQA training set, can identify words that will be used in the reference RefQ, but fail to form a coherent and answerable question. NACo, emphasizing essential criteria of a question, assigns a score of 00 to the BART-base output while giving a high score for AnnoQ ($88.89$).
Figure 4: Correlation with human judgement - Comparing CoT-QA (NACo) with Direct Evaluation
NACo vs. Existing Reference-free Metrics: Tbl. 3 also reveals that the new reference-free metric for QG, RQGUE, rates GPT-3.5 generated questions—whether in zero-shot or few-shot modes—comparably to the original reference question (RefQ). A manual review showed that GPT-3.5 typically utilizes only one of two context passages for creating a multi-hop question, as illustrated in Fig. 3. Again, human evaluation verifies our observations, detailed in Appx. A.4.2. In the case study, GPT-3.5 exclusively used context words from Passage 1, making access to Passage 2 unnecessary for answering the question. Meanwhile, RefQ incorporates context words from both passages and requires reasoning across both for an answer. RQUGE overlooks this aspect and assigns a higher score for the GPT-3.5 question than for RefQ ($94.52$ and $93.17$, respectively). Addressing this gap, NACo acknowledges the answerability and naturalness of the GPT-3.5 question, but penalizes its lower-than-expected complexity, resulting in a score of $81.48$. Since RefQ meets all three criteria of a candidate question, NACo awards it a higher score of $87.96$.
NACo (CoT-QA) vs. LLM Direct Evaluation: Large language models (LLMs) are increasingly utilized as proxies for human evaluators. We examine the effectiveness of CoT-QA used by NACo, against the conventional direct use of LLMs for evaluation. In direct evaluation, the LLMs receive the same instruction/prompt as our human evaluators. Fig. 4 presents the Pearson correlation coefficients, comparing the performance of CoT-QA (NACo) with direct evaluation across individual criteria and overall question quality. The results indicate a higher alignment with human judgment when employing CoT-QA for each respective LLM. Notably, adopting CoT-QA instead of direct evaluation significantly boosts the performance of Claude-Haiku, improving the alignment with human judgment of overall question quality from $0.46$ to $0.73$. This improvement is comparable to the performance achieved using GPT4o ($0.69$ for direct evaluation, $0.80$ in the CoTQA setting), while being $12$ times more cost effective.
@@ 5 Related Work
Evaluation Metrics for Question Generation: The evaluation of QG models commonly used reference-based metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), BERTScore Zhang et al. (2019), and BLEURT Sellam et al. (2020). Based on correlation with human judgment, there have been studies attempting to challenge the effectiveness of these reference-based metrics and propose reference-free evaluation mechanism for QG Nema and Khapra (2018); Ji et al. (2022); Mohammadshahi et al. (2023). Our study, on the other hand, questions the competency of reference-based metrics by replicating the data collection process of benchmarks and introducing new references. Other works have taken a similar approach, designing and collecting different groups of candidates to investigate reference-based metrics in machine translations Amrhein et al. (2022); Karpinska et al. (2022) and question answering Bulian et al. (2022). However, QG poses unique challenges to the evaluation of question quality, considering aspects such as complexity and answerability, and therefore call for a study like ours.
@@ 6 Conclusion
In this work, we questioned the competency of reference-based metrics in providing an accurate assessment for question generation. We replicated the data collection process used for benchmark datasets, gathering candidate questions qualified as new references. Our analysis highlights the shortcomings of reference-based metrics in differentiating new references from flawed candidates, assigning significantly lower scores to the former. Even the recently introduced reference-free metric, RQUGE, face difficulties in this regard. To address these challenges, we introduce NACo, a multi-dimensional, reference-free metric bridging the gap between automated evaluation and human judgment in question generation. Our experimental results showcase that NACo, leveraging the Chain-of-Thought capabilities of Large Language Models for question answering, not only meets the expectations for quantitative QG metrics but also achieves state-of-the-art alignment with human evaluation.
@@
A limitation of our work speaks to the required access to a reasonable number of references to assess domain-specific or dataset-specific complexity. Future works can investigate how to account for expected complexity in scenarios where references are limited and difficult to collect. Moreover, NACo, like other reference-free metrics for QG, is subject to the performance of the underlying QA model. Specifically, the constraints of GPT-3.5 in answering complex, multi-hop questions might have limited NACo’s ability to evaluate valid references closer to the upperbound. We provide a case study to illustrate this issue in Appx A.7. Future directions should explore evaluation frameworks that are robust to variations in QA model performance.
@@ A.1 What makes a good question?
Fig. 5 illustrates the gap between current automatic QG metrics and human evaluated metrics, where two questions using similar words can have opposite qualities. This gap can be explained by the fact that existing automatic metrics do not directly address any of the criteria that human annotation often looks for in a question. To address this challenge, our metric integrates the human perspective of a "good" question: naturalness, answerability, and complexity, into the evaluation pipeline.
@@ A.2 Prompt for CoT-QA
      Step 1 [reasoning step must be a single sentence with one clause]
@@
      Step 2 [reasoning step must be a single sentence with one clause]
    3. (c)
@@
A.3 NACo Details
For each question, the Chain-of-Thought (CoT) QA prompt we provide to the LLM asks the model to output its question-answering process by steps, separated by newline characters. We post-process this formatted output to count the number of reasoning steps, referred to as the absolute complexity of the candidate question or $c_{\text{cand}\textunderscore\text{abs}}$.
To calculate the relative complexity of the candidate question with respect to the dataset, we first find the expected complexity associated with that dataset. Using a set of reference questions from the training set, we perform the same CoT QA process for each of these reference questions and obtain their absolute complexity. The expected complexity for the dataset is then the most common value (or mode) among the absolute complexities of the questions in this training sample, denoted as $c_{\text{expected}}$.
The final score regarding the complexity of the candidate question is the normalized value of the absolute difference between $c_{\text{cand}\textunderscore\text{abs}}$ and $c_{\text{expected}}$: $c_{\text{cand}}=1-\frac{|c_{\text{cand}\textunderscore\text{abs}}-c_{\text{expected}}|}{\max(c_{\text{cand}\textunderscore\text{abs}},c_{\text{expected}})}$. By using $\max(c_{\text{cand}\textunderscore\text{abs}},c_{\text{expected}})$, we ensure the range of $c_{\text{cand}}$ is between 00 and $1$. The final NACo score is then computed by taking a weighted sum of $n_{\text{cand}}$ (binary, 00 or $1$), $a_{\text{cand}}$ (floating number between 00 and $1$), and $c_{\text{cand}}$ (floating number between 00 and $1$). We used a weight of $\frac{1}{3}$ for each criterion score in our experiments, ensuring NACo’s range to be between 00 and $1$. In short: $\text{NACo}=\frac{1}{3}n_{\text{cand}}+\frac{1}{3}a_{\text{cand}}+\frac{1}{3}c_{\text{cand}}$.
A.4 Human Evaluation Details
A.4.1 Instructions
@@
A.4.2 Human Evaluation Results
QG Competitor Nat. Ans. Cmp. Total
[0,2] [0,2] [0,2] [0,6]
BART-base 0.73 0.60 0.46 2.35
GPT-3.5 1.69 1.40 0.85 4.20
RefQ 1.48 1.46 1.35 4.77
AnnoQ 1.63 1.69 1.46 5.02
Table 4: Human Evaluation of QG Competitors on HotpotQA
A.5 TedEdQA Details
We collect 4246 multiple-choice questions from 1001 video lessons from TED-Ed [^2] https://ed.ted.com/. Each data point comprises the transcript of the video lesson it is based on, the question stem, and the correct answer.. After excluding questions with answers such as None of the above, All of the above, Both A and B, etc., 3547 questions remain. We split the questions into three sets train, dev, and test, each with size of 3034, 259, and 254 respectively. We ensure that no questions from any set come from the same lecture as those in the other two sets.
From the test set, we select 43 questions (RefQ) derived from 12 video lessons for additional reference annotation. We follow similar procedures to the SQuAD and HotpotQA dataset that have annotators create two types of questions—one without clues and one with provided clues—based on a given context and answer. However, the context presented to annotators differs: to formulate a reference-qualifying question, we provide them with the URL of the original lesson, the full transcript, and a specific context extracted from the transcript that is relevant to the answer. This extraction is conducted as the entire video transcript can be too long, potentially complicating the fine-tuning of models like BART. We obtain this extracted context by having GPT3.5-turbo label it from the full transcript and the original question RefQ. Specifically, we prompt the model: “Given a lecture content and a multiple-choice quiz question, please extract the most relevant and concise context from the content that is best for creating the provided multiple-choice question. Ensure the extracted excerpt contains all the necessary information for creating the given quiz question”. This context is also used to fine-tune BART-base and to generate questions with GPT-3.5-turbo in a few-shot setting.
A.6 Experiment Details
@@
For Div-Ref, which proposes diversifying references using LLM to improve reference-based metrics’ alignment with human judgement, we use the same model settings as the authors Tang et al. (2023). Specifically, we use GPT3.5-turbo with temperature set to 1 and top_p set to 0.9. We use 9/10 instructions proposed by Tang et al. 2023 to generate 9 new references from the original reference RefQ. (We did not use the remaining instruction because it asks the model to reorder sentences in a paragraph, while our text is only a question in the form of one sentence). When calculating the reference-based metric score across multiple references, we used the maximum aggregation.
LLM Details: We test 5 different LLMs for NACo. We interact with GPT3.5 (gpt3.5-turbo), GPT4o (gpt-4o) [^6] https://platform.openai.com/docs/overview, Claude-Haiku (claude-3-haiku-20240307) [^7] https://www.anthropic.com/api, and Mixtral-8x7B (open-mixtral-8x7b) [^8] https://docs.mistral.ai/api/ through their official APIs. For Llama3-8B, we download the model via their Huggingface repository (meta-llama/Meta-Llama-3-8B-Instruct) and deploy it locally. All experiments with LLMs are used with their default hyperparameters. In our CoT-QA experiments on SQuAD, given the larger sample size of 750 and cost constraints, we conducted a single run. For our CoT-QA experiments on HotpotQA, we carried out 3 runs on the 50-examples sample and report the average scores from the three responses.
A.7 Error Analysis
We have noted that one of the limitations of NACo is the dependency of the QA models’ performance. To further elaborate it, we provide the a case study in Fig. 6. The case study involves two context passages and the answer ‘Teinosuke Kinugasa.’ We examine three candidate questions: a GPT-3.5-generated question (intended for 2-hop but resulting in 1-hop), the original HotpotQA reference (RefQ, 2-hop), and our newly collected reference (AnnoQ-clue-RefQ, 2-hop). The CoT-QA model employed by NACo (GPT3.5) correctly identifies ’Teinosuke Kinugasa’ for both the GPT-3.5-generated question and AnnoQ-clue-RefQ but fails to do so for RefQ, responding with ‘Not enough information provided to answer the question.’
This failure with RefQ is attributed to its requirement for mathematical reasoning (subtracting birth year from death year), a task GPT-3.5 struggles with. Accordingly, NACo assigns the highest score to AnnoQ-clue-RefQ, fulfilling all three requirements, while penalizing the GPT-3.5 question for its simplicity and RefQ most severely (Answerability F1 score = 0) due to the mismatch between the CoT-QA answer and the provided answer.
Figure 6: NACo Error Analysis: Reliance on QA model capacity.
A.8 Additional results for reference-based metrics
Tbl. 5 and 6 provides a more detailed version of 1 and 3.
Tbl. 7 illustrates the application of our study that disproves reference-based metrics and the proposed metric, NACo, in an educational setting using the TedEd-QA dataset. It can be seen that our observations regarding the failure of reference-based metrics and the effectiveness of NACo also holds for this dataset. Specifically, Refq, AnnoQ, and AnnoQ-clue-RefQ have significant gap when reference-based metrics are used to score them. NACo is able to score all these three human-validated candidates with similar scores and no worse than any machine-generated candidates.
QG Competitor Ref-based metrics Ref-free metrics
B B-RT R-L BSc Q-B QA-S R-Q NACo
LM-generated
BART-base 19.53 -0.28 44.79 92.13 36.94 -0.37 4.62 73.30
GPT-3.5 (few-shot) 18.06 -0.23 43.58 92.18 36.48 -0.37 4.56 73.67
BART-clue-RefQ 31.91 0.07 59.92 94.37 52.33 -0.38 4.56 69.97
Human-validated
RefQ 100.00 1.00 100.00 100.00 100.00 -0.38 4.89 75.09
AnnoQ 12.78 -0.31 37.83 91.52 31.32 -0.37 4.71 74.01
AnnoQ-clue-RefQ 27.43 0.04 53.62 93.85 46.89 -0.38 4.76 74.21
Table 5: SQuAD - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original SQuAD questions (RefQ) as references, with GPT-3.5 as the underlying QA system for NACo. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
QG Competitor Ref-based metrics Ref-free metrics Human
B B-RT R-L BSc Q-B QA-S R-Q NACo
LM-generated
BART-base 12.97 -0.79 34.38 87.88 27.67 -0.28 2.78 41.25 2.35
GPT-3.5 (zero-shot) 7.94 -0.95 26.05 87.38 17.00 -0.28 4.25 73.60 4.20
GPT-3.5 (few-shot) 10.53 -0.87 27.51 87.76 18.67 -0.27 4.21 71.09 -
BART-clue-RefQ 31.90 -0.10 58.77 92.61 51.98 -0.28 3.31 63.17 -
Human-validated
RefQ 100.00 1.00 100.00 100.00 100.00 -0.27 4.21 74.33 4.77
AnnoQ 13.64 -0.70 30.94 89.22 26.95 -0.27 4.32 83.43 5.02
AnnoQ-clue-RefQ 35.59 -0.06 52.95 92.15 51.76 -0.27 4.44 81.60 -
Table 6: HotpotQA - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original HotpotQA questions (RefQ) as references, with GPT-3.5 as the underlying QA system for NACo. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
QG Competitor Ref-based metrics Ref-free metrics Our metric
B B-RT R-L BSc Q-B QA-S R-Q NACo
LM-generated
BART-base 13.71 0.16 32.42 89.19 - -0.15 3.74 79.29
GPT-3.5 (few-shot) 9.23 -0.28 28.77 88.6 - -0.16 3.89 79.78
BART-clue-RefQ 13.63 0.01 40.66 90.09 - -0.15 3.33 75.29
Human-validated
RefQ 100.00 1.00 100.00 100.00 - -0.16 3.99 82.85
AnnoQ 14.78 -0.43 34.16 89.61 - -0.16 3.98 84.67
AnnoQ-clue-RefQ 26.4 0.59 50.22 92.1 - -0.17 4.23 83.00
Table 7: TedEdQA - Performance of different QG methods using NACo and other existing metrics. The evaluation uses original TedEd questions (RefQ) as references, with GPT-3.5 as the underlying QA system for NACo. Some questions in this dataset are in the fill-in-the-blank form and do not contain question words like what, where, etc. which Q-BLEU heavily relies on for scoring Nema and Khapra (2018); thus, we do not report this metric for this dataset. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.